75 g of water at  69 ∘C are mixed with 75 g of water at 20 ∘C in a thermally-insulated container.

Calculate the change in entropy, ΔS.

I know the answer is  ΔS = 1.87 J/K, but I cannot remember what I did to get this last week. Thank you!
Answer #1

Let Tf be the final temperature. Mass of each water = 75 g heat energy associated with hot water is hot = m* C* AT = 75 g * 4

Tf = 44.5 °C = 44.5 + 273.15 = 317.65 K

Hot temperature = 69 °C = 69 + 273.15 = 342.15 K

Cold temperature = 20 °C = 20 + 273.15 = 293.15 K

Specific heat capacity = 4.184 J/(°C.g) = 4.184 J/(K.g)

AS = .x XComax on it and Co k tomu na 317.6 317.65 S = 75 g x 4.184 J/(K.g) x In + 75 gx 4.184 J/(°C.g)x In – 293.15) 317.65